### v4

@@ 2.1 The Extended Robust UCB Policy
Consider $K$ arms offering random rewards $\{X_{1},X_{2},\dots,X_{K}\}$ with heavy-tailed distributions =$\mathcal{F}=\{F_{1},F_{2},\dots,F_{K}\}$. In contrast with the light-tailed class, a heavy-tailed distribution may yield a very high (or low) reward realization with a high probability. Consequently, the empirical mean estimator $\overline{X}_{i,s}=\frac{1}{s}\sum_{j=1}^{s}X_{i,j}$ for a sequence of i.i.d. random variables $\{X_{i,j}\}_{j=1}^{s}$ may not work properly to estimate the expectation of $X_{i}$ for arm $i$. The appendix of  Bubeck et al. [5] has shown that under the assumption of heavy-tailed distributions, with probability at least $1-\delta$,
@@
where $1<r\leq 2$. This bound is too loose to obtain a UCB policy for achieving logarithmic regret growth. To remedy this issue, we adopt the median-of-means estimator proposed in  Bubeck et al. [5].
@@
$\mathbb{E}[|X_{i}-\theta_{i}|^{p}]\leq C_{p,q}\mathbb{E}[|X_{i}-\theta_{i}|^{q}]^{r},$ (6)
@@
Let $k=\lceil 8\log\delta^{-1}\rceil,\chi=\frac{(8+\epsilon)\log\delta^{-1}}{s}$. Given an arbitrary $\epsilon>0$, a moment order $p>1$, another moment order $1<q<p$, and the moment control coefficient $C_{p,q}$, define the upper confidence bound of the extended robust UCB as (value of the fraction is set to +$\infty$+ ∞ if divided by 00)
@@
Note that this upper confidence bound always exists, since the left-hand side of the above inequality tends to $\infty$- ∞ as $\theta\to-\infty$, while the right-hand side tends to +$\infty$+ ∞.
@@
                  Set arm $i$’s upper confidence bound to +$\infty$+ ∞
@@
Suppose $p>q>1$ such that the $p$-th order moments exist for all $F_{i}\in\mathcal{F}$ and (6) is satisfied with known $C_{p,q}$. Then for any $\epsilon>0$, the regret $\mathcal{R}_{\pi^{*}}^{\mathcal{F}}(T)$ of the extended robust UCB policy $\pi^{*}$ has the logarithmic order with respect to the time horizon $T$.
To prove Theorem 1, we just need to show the following core lemma to bound the time $s_{i}$ spent on each suboptimal arm $i$ by the extended robust UCB policy.
@@
The expected number of times $\mathbb{E}[s_{i}(T)]$ that a suboptimal arm is chosen by the extended robust UCB policy $\pi^{*}$ has an upper bound of the logarithmic order with $T$.
A Monte Carlo simulation example based on the Pareto distribution of type III (see Sec. 1.1) is shown in Fig. 1. Note that the regret rate grows linearly at the beginning to prepare the ground for ranking until at least one arm has a finite UCB (i.e. $\chi$ defined in (24) becomes sufficiently small such that, within the $\sup$ operator of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}},t^{-2})$, both the denominator is not zero and the growth rate of the right-hand side with $\theta$ becomes small). Note that if $C_{p,q}$ is unknown, then the player can simply choose any increasing function $f$ with $f(t)\rightarrow+\infty$ as $t\rightarrow+\infty$ to replace $C_{p,q}$ and achieve a regret arbitrarily close to logarithmic order without requiring any knowledge of the reward moments (see Appendix F for a proof).
Figure 1: The performance of extended robust UCB
@@ 2.2 Further Analysis and Comparison of The Extended Robust UCB
Now we consider scenarios of specific chosen $p$ and $q$ to further improve the efficiency of the extended robust UCB policy.
Case 1. Suppose that the $p^{\prime}$-th order of moments of reward distributions exists with $p^{\prime}>2$ for all arms. Then we can choose $p$ and $q$ such that $p=2q\leq p^{\prime}$. The inequality (29) can be refined to
@@
Now we look at the definition of $\chi_{M}$ in Appendix E that ensures the three inequalities (39), (40) and (41) to hold. The refinement of $C^{\prime}$ leads to a smaller right-hand side for all these inequalities, i.e. the constant $\chi_{M}$ will be larger and the regret upper bound will be smaller, which further implies that the policy’s selection will converge to the best arm faster.
@@
Under these refined estimations for $\chi_{M}$, the leading coefficient of $\log T$ in (46) is significantly decreased. Now we compare the extended robust UCB and the scale free algorithm proposed in  Lattimore [22], under the same assumption specified at the beginning of Case 2.
@@
A Flawed Statement  Bubeck et al. [5]: Let $\delta\in(0,1)$ and $p\in(1,2]$. Let $\{X_{i,j}\}_{j=1}^{s}$ be i.i.d. random variables with mean $\mathbb{E}X_{i}=\theta_{i}$ and (centered) $p$-th moment $v_{p}$. Let $k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor$. Then with probability at least $1-\delta$,
@@
Last, we show that the regret of the extended robust UCB deviates from the theoretical lower bound (in the non-parametric setting) by only a constant factor and a constant term as in the scale free algorithm of  Lattimore [22]. First, we restate the lower bound derived in  Lattimore [22] as follows.
@@
[…]
@@ 3 Light-Tailed Reward Distributions
We point out that our policy and results also work for the light-tailed distributions since our construction depends only on the existence of moments. However, the leading constant of the logarithmic regret order can be improved by our past work in Liu and Zhao [24], which we formally and briefly discuss here with a comparison with some subsequent work following Liu and Zhao [24].
@@ 3.1 The UCB1-LT Policy
                  Assign +$\infty$+ ∞ to the upper confidence bound
@@ Appendix A From UCB1-LT to $(\alpha,\psi)$-UCB
$\log\mathbb{E}\left[{\rm e}^{\lambda(X_{i}-\theta_{i})}\right]\leq\psi(\lambda)\text{ and }\log\mathbb{E}\left[{\rm e}^{\lambda(\theta_{i}-X_{i})}\right]\leq\psi(\lambda).$ (17)
For example, if $X_{i}$ is bounded in $[0,1]$ as assumed in UCB1, one can choose $\psi(\lambda)=\frac{\lambda^{2}}{8}$ and (17) becomes the well-known Hoeffding’s lemma [4]. Since +$\infty$+ ∞ is allowed in the range of $\psi$, this assumption is more general than that in UCB1-LT. The Legendre-Frenchel transform $\psi^{*}$ of $\psi$ is defined as
@@
Now we compare the actual performance between $(\alpha,\psi)$-UCB and UCB1-LT through numerical experiments. Assume that the arms are offering rewards according to $19$ normal distributions. The means of reward distributions are $0.5,1.0,1.5,\dots,9.5$ with standard deviations all set to $60$. A Monte Carlo (MC) simulation with 100 runs and time period $T=50000$ is shown below. The horizontal axis in both figures is the time step $t$. The first figure only shows $T$ from $10000$ to $50000$ for better display, while the second one ranges from $1$ up to $50000$. The vertical axis of the left figure is the logarithmic regret averaged from the 100 MC runs (to approximate the expectation), while of the right one is the approximately expected time-average reward. The parameters for UCB1-LT and $(\alpha,\psi)$-UCB are respectively $u_{0}=1$ and $\alpha=2.5$ or $4$ with $\zeta=3600$ as their common parameter.
@@ Appendix B Proof of Lemma 1
Note that $N=\lfloor s/\lceil 8\log\delta^{-1}\rceil\rfloor$. Since $s\geq(8+\epsilon)\log\delta^{-1}\frac{8\log(\delta^{-1})+1}{\epsilon\log(\delta^{-1})-1}$, a direct computation yields $\xi\leq 1/4$. Then using Hoeffding’s inequality for the tail of a binomial distribution Hoeffding [15], we have
@@
Similarly, the other inequality also holds by symmetry. $\hfill\qed$
@@ Appendix C Proof of Lemma 2
The proof is thus completed by the definition of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)$. $\hfill\qed$
@@ Appendix D Proof of Corollary  1
This corollary is a direct consequence of Lemma 1 and the inequality in (29). $\hfill\qed$
@@ Appendix E Proof of Lemma 3
by the convexity of $|x|^{q}$ ($q>1$) and Jensen’s inequality Rudin [29]. Suppose that
@@
[…]
@@
We will bound the first part of the right hand side of (LABEL:s_i_divide) by Lemma 2. Note that $\log\delta^{-1}>1/\epsilon$ is equivalent to $t>\exp\left(\frac{1}{2\epsilon}\right)$. Write
@@
Also note that for sufficiently large $T$, the function $\log x^{2}\frac{8\log x^{2}+1}{\epsilon\log x^{2}-1}$ is increasing for $x\in[l_{0}+1,+\infty)$. We thus have
@@
Now we bound the second part of the right hand side of (LABEL:s_i_divide). Define
@@
For sufficiently large $T$, we can assume that $l_{0}=\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil$. Thus the above inequality becomes
$\begin{split}\mathbb{E}[s_{i}(T)]&\leq\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil\\ &+\max\left\{\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil,\left\lceil\frac{8+\epsilon}{\chi_{M}}\log T^{2}\right\rceil\right\}+\frac{5\pi^{2}}{6}.\end{split}$ (46)
$\hfill\qed$

### v5

@@ 2.1 The Extended Robust UCB Policy
Consider $K$ arms offering random rewards $\{X_{1},X_{2},\dots,X_{K}\}$ with heavy-tailed distributions $\mathcal{F}=\{F_{1},F_{2},\dots,F_{K}\}$. In contrast with the light-tailed class, a heavy-tailed distribution may yield a very high (or low) reward realization with a high probability. Consequently, the empirical mean estimator $\overline{X}_{i,s}=\frac{1}{s}\sum_{j=1}^{s}X_{i,j}$ for a sequence of i.i.d. random variables $\{X_{i,j}\}_{j=1}^{s}$ may not work properly to estimate the expectation of $X_{i}$ for arm $i$. The appendix of Bubeck et al. [5] has shown that under the assumption of heavy-tailed distributions, with probability at least $1-\delta$,
@@
where $1<r\leq 2$. However, this bound is too loose to obtain a UCB policy for achieving logarithmic regret growth. To remedy this issue, we adopt the median-of-means estimator proposed in Bubeck et al. [5].
@@
$\mathbb{E}[|X_{i}-\theta_{i}|^{p}]\leq C_{p,q}[\mathbb{E}[|X_{i}-\theta_{i}|^{q}]]^{r},$ (6)
@@
Let $k=\lceil 8\log\delta^{-1}\rceil,\chi=\frac{(8+\epsilon)\log\delta^{-1}}{s}$. Given an arbitrary $\epsilon>0$, a moment order $p>1$, another moment order $1<q<p$, and the moment control coefficient $C_{p,q}$, define the upper confidence bound of the extended robust UCB as (value of the fraction is set to $\infty$ if divided by 0)
@@
Note that this upper confidence bound always exists, since the left-hand side of the above inequality tends to -$\infty$ as $\theta\to$-$\infty$, while the right-hand side tends to $\infty$.
@@
        Set arm $i$’s upper confidence bound to $\infty$
@@
Suppose $p>q>1$ such that the $p$-th order moments exist for all $F_{i}\in\mathcal{F}$ and (6) is satisfied with known $C_{p,q}$. Then for any $\epsilon>0$, the regret $\mathcal{R}_{\pi^{*}}^{\mathcal{F}}(T)$ of the extended robust UCB policy $\pi^{*}$ has the logarithmic order with respect to the time horizon $T$. Specifically, we have $\mathcal{R}_{\pi^{*}}^{\mathcal{F}}(T)\leq\sum_{i:\Delta_{i}>0}C_{i}(T)\Delta_{i}$ with
$C_{i}(T):=\Big{[}36(1+\frac{8}{\epsilon})+(16+2\epsilon)\Big{(}144(1+C_{p,q})(6v_{q}^{\frac{1}{q}}+2^{q})(\Delta_{i}+\frac{1}{\Delta_{i}})\Big{)}^{M(p,q)}\Big{]}\log T+11,$
where $M(p,q):=1/((1-\frac{1}{q})(1-\frac{q}{p}))$ and $T\geq\frac{9}{2\epsilon}+\frac{3}{2}.$
From the above, we observe that the upper bound on regret increases with $C_{p,q}$ but decreases with $p$. This matches the intuition that a tighter known bound on the moment ratio or a higher moment order adopted makes the learning process easier due to less variation assumed in the samples. Furthermore, for any fixed $q$, the upper bound increases with $v_{q}$. This is also easy to explain since a larger moment indicates a larger deviation from a sample to the true mean. Meanwhile, if $q$ increases and becomes close to $p$, then the denominator of $M(p,q)$ will approach 0 and the upper bound becomes worse. This is because a larger value of $p/q$ offers more global information on moments with different orders to facilitate the learning of the unknown distributions. Also, as $\Delta_{i}\rightarrow 0$ for some $i$, the upper bound goes to infinity (since $M(p,q)>1$) because it will be harder to distinguish the best arm from arm $i$. Finally, if the arm number $K$ is increased with other parameters including $\Delta_{i}$ relatively fixed within the original orders, the upper bound is clearly increased linearly with $K$ due to the definition of regret. To prove Theorem 1, we just need to show the following core lemma to bound the time $s_{i}$ spent on each suboptimal arm $i$ by the extended robust UCB policy.
@@
The expected number of times $\mathbb{E}[s_{i}(T)]$ that a suboptimal arm is chosen by the extended robust UCB policy $\pi^{*}$ has an upper bound of the logarithmic order with $T$. Specifically, we have $\mathbb{E}[s_{i}(T)]\leq C_{i}(T)$ where $C_{i}(T)$ is defined above.
Finally, we conduct Monte Carlo simulation examples based on Laplace distributions, Student-t distributions, and Pareto distributions of type III (see Sec. 1.1). From Fig. 14, we observe that the regret rate grows linearly at the beginning to prepare the ground for ranking until at least one arm has a finite UCB (i.e., $\chi$ defined in (24) becomes sufficiently small such that, within the $\sup$ operator of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}},t^{-2})$, not only the denominator is nonzero but also the growth rate of the right-hand side with $\theta$ becomes small). In Fig. 1, we compare the actual regret for different values of $C_{p,q}$ and observe that the regret becomes larger for larger $C_{p,q}$. In Fig 2, the regret becomes smaller as $p$ becomes larger. Furthermore, the regret increases with $v_{q}$ and $K$ as shown in Fig. 3 and Fig. 4, respectively. All these observations match our theoretical analysis on the upper bound above.
(a) Laplace with $C_{p,q}=1.68\times 10^{3}$
(b) Student-t with $C_{p,q}=2.09\times 10^{3}$
Figure 1: Comparison for different $C_{p,q}$
(a) Pareto III with $p=12$
(b) Pareto III with $p=15$
Figure 2: Comparison for different $p$
(a) Laplace with $v_{q}=24$
(b) Laplace with $v_{q}=384$
Figure 3: Comparison for different $v_{q}$
(a) Student-t with $K=2$
(b) Student-t with $K=4$
Figure 4: Comparison for different $K$
Note that if $C_{p,q}$ is unknown, then the player can simply choose any increasing function $f$ with $f(t)\rightarrow+\infty$ as $t\rightarrow+\infty$ to replace $C_{p,q}$ and achieve a regret arbitrarily close to logarithmic order without requiring any knowledge of the reward moments (see Appendix F for a proof).
@@ 2.2 Further Analysis and Comparison of The Extended Robust UCB
Now we consider the specific chosen values of $p$ and $q$ to further improve the efficiency of the extended robust UCB policy.
Case 1. Suppose that the $p^{\prime}$-th order of moments of reward distributions exists with $p^{\prime}>2$ for all arms. Then we can choose $p$ and $q$ such that $p=2q\leq p^{\prime}$. Inequality (29) can be refined to
@@
Next, we look at the definition of $\chi_{M}$ in Appendix E that guarantees the three inequalities (39), (40) and (41) to hold. The refinement of $C^{\prime}$ leads to a smaller right-hand side for all these inequalities, i.e. the constant $\chi_{M}$ will be larger and the regret upper bound will be smaller, which further implies that the policy’s selection will converge to the best arm faster.
@@
Under these refined estimations for $\chi_{M}$, the leading coefficient of $\log T$ in (46) is significantly reduced. Now we compare the extended robust UCB and the scale free algorithm proposed in Lattimore [22], under the same assumption specified at the beginning of Case 2.
@@
A Flawed Statement in Bubeck et al. [5]: Let $\delta\in(0,1)$ and $p\in(1,2]$. Let $\{X_{i,j}\}_{j=1}^{s}$ be i.i.d. random variables with mean $\mathbb{E}X_{i}=\theta_{i}$ and (centered) $p$-th moment $v_{p}$. Let $k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor$. Then with probability at least $1-\delta$,
@@
Furthermore, we show that the regret of the extended robust UCB deviates from the theoretical lower bound (in the non-parametric setting) by only a constant factor and a constant term as in the scale free algorithm of Lattimore [22]. First, we restate the lower bound derived in Lattimore [22] as follows.
@@
where $C_{0}$, $C_{1}>0$ are universal constants and $p=\min\{\Delta/\sigma,\ 1/\kappa_{0}\}$.
@@
Furthermore, we give another upper bound that shows a sublinear regret in terms of arm number $K$. Recall the upper bound on regret in Theorem 1. We have
$\mathcal{R}_{\pi^{*}}^{\mathcal{F}}(T)$ $=\sum_{i:\Delta_{i}>0}\Delta_{i}\mathbb{E}[s_{i}(T)]=\sum_{i:\Delta_{i}>0}\Delta_{i}[\mathbb{E}[s_{i}(T)]]^{1/M(p,q)}[\mathbb{E}[s_{i}(T)]]^{1-1/M(p,q)}$
$\leq\sum_{i:\Delta_{i}>0}\Delta_{i}\Big{[}\Big{(}M_{1}+M_{2}(p,q)(\frac{\Delta_{i}^{2}+1}{\Delta_{i}})^{M(p,q)}\Big{)}2\log T\Big{]}^{1/M(p,q)}[\mathbb{E}[s_{i}(T)]]^{1-1/M(p,q)}$
$\leq\sum_{i:\Delta_{i}>0}M_{3}(p,q)(\Delta_{i}^{2}+1)(\log T)^{1/M(p,q)}[\mathbb{E}[s_{i}(T)]]^{1-1/M(p,q)}$
$\leq M_{3}(p,q)(\log T)^{1/M(p,q)}\Big{[}\sum_{i:1\geq\Delta_{i}>0}2[\mathbb{E}[s_{i}(T)]]^{1-1/M(p,q)}+$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\sum_{i:\Delta_{i}>1}2\Delta_{i}^{2}[\mathbb{E}[s_{i}(T)]]^{1-1/M(p,q)}\Big{]}$
$\leq 2M_{3}(p,q)(\log T)^{1/M(p,q)}\Big{[}K^{1/M(p,q)}T^{1-1/M(p,q)}+\sum_{i:\Delta_{i}>1}\Delta_{i}^{2}T^{1-1/M(p,q)}\Big{]}$
$\leq M_{3}(p,q)(\log T)^{1/M(p,q)}T^{1-1/M(p,q)}\Big{[}K^{1/M(p,q)}+\sum_{i:\Delta_{i}>1}\Delta_{i}^{2}\Big{]}$
where $M_{1}=19+144/\epsilon,~M(p,q)=1/((1-1/q)(1-q/p))>1,~T\geq\frac{9}{2\epsilon}+\frac{3}{2}$, and
$M_{2}(p,q)$ $=(8+\epsilon)\Big{(}144(1+C_{p,q})(6v_{q}^{\frac{1}{q}}+2^{q})\Big{)}^{M(p,q)},$
$M_{3}(p,q)$ $=\Big{(}M_{1}+2M_{2}(p,q)\Big{)}^{1/M(p,q)}.$
It is desirable to remove the dependency of regret bound on $\Delta_{i}$ as well and this will be considered in the future work.
@@ 3 Light-Tailed Reward Distributions
We point out that our policy and results also work for the light-tailed reward distributions since our construction depends only on the existence of moments. However, as mentioned in Sec. 1.2, the UCB could be significantly simplified for the light-tailed class where learning reward means becomes much easier. Furthermore, the leading constant of the logarithmic regret upper bound can also be greatly reduced as shown in [24]. Therefore, we now formally but briefly discuss the work in [24] that extended the UCB policy for distributions with known finite supports in [3] to all light-tailed distributions. We will also compare our work with some subsequent work following [24]. Together with above results for the heavy-tailed class, we see a complete picture of the beauty and power of UCB as well as what information needs to be extracted in its specific design based on how the samples represent the true means.
@@ 3.1 The UCB1-LT Policy
        Assign $\infty$ to the upper confidence bound
@@ Appendix A From UCB1-LT to $(\alpha,\psi)$-UCB
[…]
For example, if $X_{i}$ is bounded in $[0,1]$ as assumed in UCB1, one can choose $\psi(\lambda)=\frac{\lambda^{2}}{8}$ and (17) becomes the well-known Hoeffding’s lemma [4]. Since $\infty$ is allowed in the range of $\psi$, this assumption is more general than that in UCB1-LT. The Legendre-Frenchel transform $\psi^{*}$ of $\psi$ is defined as
@@
Now we compare the actual performance between $(\alpha,\psi)$-UCB and UCB1-LT through numerical experiments shown in Fig. 5. Assume that the arms are offering rewards according to $19$ normal distributions. The means of reward distributions are $0.5,1.0,1.5,\dots,9.5$ with standard deviations all set to $60$. A Monte Carlo (MC) simulation with 100 runs and time period $T=50000$ is shown below. The horizontal axis in both figures is the time step $t$. The first figure only shows $T$ from $10000$ to $50000$ for better display, while the second one ranges from $1$ up to $50000$. The vertical axis of the left figure is the logarithmic regret averaged from the 100 MC runs (to approximate the expectation), while of the right one is the approximately expected time-average reward. The parameters for UCB1-LT and $(\alpha,\psi)$-UCB are respectively $u_{0}=1$ and $\alpha=2.5$ or $4$ with $\zeta=3600$ as their common parameter.
@@ Appendix B Proof of Lemma 1
Note that $N=\lfloor s/\lceil 8\log\delta^{-1}\rceil\rfloor$. Since $s\geq(8+\epsilon)\log\delta^{-1}\frac{8\log(\delta^{-1})+1}{\epsilon\log(\delta^{-1})-1}$, a direct computation yields $\xi\leq 1/4$. Then using Hoeffding’s inequality for the tail of a binomial distribution [15], we have
@@
[…]
@@ Appendix C Proof of Lemma 2
[…]
@@ Appendix D Proof of Corollary  1
[…]
@@ Appendix E Proof of Lemma 3
by the convexity of $|x|^{q}$ ($q>1$) and Jensen’s inequality [29]. Suppose that
@@
Observe that both $\tau(x)$ andB(x)B(x) decrease as $x$ decreases and
@@
We will bound the first part of the right-hand side of (LABEL:s_i_divide) by Lemma 2. Note that $\log\delta^{-1}>1/\epsilon$ is equivalent to $t>\exp\left(\frac{1}{2\epsilon}\right)$. Write
@@
Also note that for sufficiently large $T$, the function $\log x^{2}\frac{8\log x^{2}+1}{\epsilon\log x^{2}-1}$ is increasing for $x\in[l_{0}+1,+\infty)$. We then have
@@
Now we bound the second part of the right-hand side of (LABEL:s_i_divide). Define
@@
For sufficiently large $T$, we can assume that $l_{0}=\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil$. Therefore, the above inequality becomes
$\begin{split}\mathbb{E}[s_{i}(T)]&\leq\left\lceil(8+\epsilon)\log(T^{2})\frac{8\log(T^{2})+1}{\epsilon\log(T^{2})-1}\right\rceil\\ &+\max\left\{\left\lceil(8+\epsilon)\log(T^{2})\frac{8\log(T^{2})+1}{\epsilon\log(T^{2})-1}\right\rceil,\left\lceil\frac{8+\epsilon}{\chi_{M}}\log(T^{2})\right\rceil\right\}+\frac{5\pi^{2}}{6}.\end{split}$ (46)
[…]